# Availability-aware Sensor Fusion via Unified Canonical Space for 4D Radar, LiDAR, and Camera

### v1

@@ 3.2 Availability-aware Sensor Fusion (ASF)
where $C_{q}$ is the quotient of $C_{u}/(P_{H}\!\times\!P_{W})$ as we design $C_{u}\!=\!P_{H}\!\times\!P_{W}\!\times\!C_{q}$. Since the resulting channel dimension $C_{q}$ may be insufficient for containing feature representation due to reduced channel dimensions after reshaping, in the implementation, we increase the number of patches by a factor of $n_{p}$ (i.e., $N_{p}\!=\!N_{H}\!\times\!N_{W}\!\rightarrow\!N_{p}\!=\!n_{p}\!\times\!N_{H}\!\times\!N_{W}$). Consequently, the channel dimension of $\mathbf{FM}_{\mathrm{fused}}$ increases from $C_{q}$ to $n_{p}\!\times\!C_{q}$, and the impact of this modification is evaluated in subsection 4.4.
Methods Sensors IoU Metric Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
RTNH [12] (NeurIPS 2022) R 0.3 $AP_{BEV}$ 41.1 41.0 44.6 45.4 32.9 50.6 81.5 56.3
$AP_{3D}$ 37.4 37.6 42.0 41.2 29.2 49.1 63.9 43.1
0.5 $AP_{BEV}$ 36.0 35.8 41.9 44.8 30.2 34.5 63.9 55.1
$AP_{3D}$ 14.1 19.7 20.5 15.9 13.0 13.5 21.0 6.36
RTNH [12] (NeurIPS 2022) L 0.3 $AP_{BEV}$ 76.5 76.5 88.2 86.3 77.3 55.3 81.1 59.5
$AP_{3D}$ 72.7 73.1 76.5 84.8 64.5 53.4 80.3 52.9
0.5 $AP_{BEV}$ 66.3 65.4 87.4 83.8 73.7 48.8 78.5 48.1
$AP_{3D}$ 37.8 39.8 46.3 59.8 28.2 31.4 50.7 24.6
3D-LRF [2] (CVPR 2024) L+R 0.3 $AP_{BEV}$ 84.0 83.7 89.2 95.4 78.3 60.7 88.9 74.9
$AP_{3D}$ 74.8 81.2 87.2 86.1 73.8 49.5 87.9 67.2
0.5 $AP_{BEV}$ 73.6 72.3 88.4 86.6 76.6 47.5 79.6 64.1
$AP_{3D}$ 45.2 45.3 55.8 51.8 38.3 23.4 60.2 36.9
L4DR [34] (AAAI 2025) L+R 0.3 $AP_{BEV}$ 79.5 86.0 89.6 89.9 81.1 62.3 89.1 61.3
$AP_{3D}$ 78.0 77.7 80.0 88.6 79.2 60.1 78.9 51.9
0.5 $AP_{BEV}$ 77.5 76.8 88.6 89.7 78.2 59.3 80.9 53.8
$AP_{3D}$ 53.5 53.0 64.1 73.2 53.8 46.2 52.4 37.0
ASF (Proposed) L+R 0.3 $AP_{BEV}$ 88.6 88.1 90.3 99.0 89.1 80.4 89.4 78.7
$AP_{3D}$ 87.3 86.6 89.8 90.7 88.6 80.0 88.8 77.5
0.5 $AP_{BEV}$ 87.0 86.2 90.2 90.8 88.8 78.2 88.6 71.0
$AP_{3D}$ 72.9 64.6 86.6 79.6 73.4 70.0 77.6 66.7
C+L+R 0.3 $AP_{BEV}$ 88.6 88.2 90.2 98.9 89.0 80.4 89.2 78.4
$AP_{3D}$ 87.4 87.0 90.1 90.7 88.2 80.0 88.6 77.4
0.5 $AP_{BEV}$ 87.2 86.7 90.1 90.8 88.7 78.3 88.3 70.9
$AP_{3D}$ 73.6 71.8 87.0 79.4 73.0 67.5 78.0 66.4
Table 1: Performance comparison of 3D object detection on K-Radar [12] benchmark v1.0. C, L, and R represent Camera, LiDAR, and 4D Radar, respectively. The bold and underlined values indicate the best and the second-best, respectively. Note that the two ASF models for sensors L+R and C+L+R share the same neural network weights trained with C+L+R, which means that ASF for L+R represents a scenario where camera becomes unavailable.
@@ 3.3 Sensor Combination Loss (SCL)
where $\mathcal{SC}$ represents the set of $7$ possible sensor combinations ($S_{C}$-only, $S_{L}$-only, $S_{R}$-only, $S_{L}\!+\!S_{R}$, $S_{C}\!+\!S_{R}$, $S_{C}\!+\!S_{L}$, and $S_{C}\!+\!S_{L}\!+\!S_{R}$), where $S_{C}$, $S_{L}$, and $S_{R}$ denote camera, LiDAR, and 4D Radar, respectively. $\mathcal{L}_{cls,s}$ and $\mathcal{L}_{reg,s}$ are the classification and regression losses for each sensor combination. SCL explicitly prepares for the potential sensor unavailability by optimizing across all sensor combinations in the training, enabling the model to recognize that certain sensors perform better than others (e.g., 4D Radar and LiDAR outperform cameras in adverse weather). As demonstrated in Tab. 4, SCL enhances the performance of the proposed ASF method when compared to ASF without SCL.
@@ 4.1 Experimental Setup
Dataset and Metrics. K-Radar [12] is a large-scale AD dataset with a broad range of conditions including time (day, night), weather (normal, rain, fog, snow, sleet), road types (urban, highway, mountain), and sensors (4D Radar, LiDAR, camera, GPS). Notably, K-Radar is the only dataset with data captured in adverse weather conditions.
To compare ASF with the existing SOTA methods, we use two official variants of K-Radar settings. Following the protocol in [12] and [2, 34], we evaluate ASF and SOTA methods using benchmark v1.0, which focuses on the ‘Sedan’ class with label v1.0 and covers a region close to the driving corridor: [0m, 72m] for X-axis, [-6.4m, 6.4m] for Y-axis, and [-2m, 6m] for Z-axis. Additionally, we present ablation studies and qualitative results on benchmark v2.0, which analyzes a wider area ([0m, 72m] for X-axis, [-16m, 16m] for Y-axis, and [-2m, 7.6m] for Z-axis) and includes both ‘Sedan’ and ‘Bus or Truck’ classes (two dominant classes in [12]). For evaluation metrics, we use: $AP_{3D}$ and $AP_{BEV}$ at IoU thresholds of $0.3$ (following the protocol in [12]) and $0.5$ (for a strict assessment). Additionally, we report the required VRAM and FPS (frames per second) based on the same hardware setup.
Implementation Details. We implement the ASF on a single RTX3090 GPU with 24GB VRAM. ASF is trained for 11 epochs using AdamW [44] optimizer with a learning rate $0.001$ and a batch size $2$. The voxel size for the fused FM is set to $0.4$m, consistent with [12].
Figure 3: Qualitative results of ASF for various sensor combinations. We show results for normal and adverse weather conditions in (a-i) and (j-r), respectively, where employed sensors are noted in the top-left corner (C: Camera, L: LiDAR, R: 4D Radar, C*: damaged camera). Each subplot visualizes front-view camera image, LiDAR point cloud, 4D Radar tensor, and a sensor attention map (SAM) showing attention score distribution from cross-attention in CASAP-PN. In the SAMs, red, green, and blue represent attention scores for Camera, LiDAR, and 4D Radar, respectively. For example, a predominantly blue SAM indicates that 4D Radar received the highest attention scores in the scene. The bottom-left corner of each subplot shows the proportion of attention scores in colored percentages (C/L/R[%]). Note that predictions are visualized on all sensor data, even when a sensor is not employed for detection (e.g., predictions from L+R are also visualized on the camera image).
4.2 Comparison of ASF to SOTA methods
Following the benchmark v1.0 of the K-Radar [12], we compare the proposed ASF with SOTA methods including 3D-LRF [2] and L4DR [34], and we use RTNH [12] for single-sensor performance. In addition to the detection performance comparison with DCF methods [2, 34], we evaluate computational efficiency against DPFT [3] which is the only open-sourced SCF method available for K-Radar.
Methods Sensors VRAM [GB] FPS [Hz]
3D-LRF [2] L+R 1.2 5.04
DPFT [3] C+R 4.0 10.8
ASF L+R 1.5 20.5
(Proposed) C+L+R 1.6 13.5
Table 2: Comparison of VRAM and FPS evaluated on the K-Radar benchmark v1.0. The bold and underlined values indicate the best and the second-best, respectively.
Class Sensors Total Normal Ove. Sleet H.s.
@@ 4.2 Comparison of ASF to SOTA methods
Table 3: Performance comparison of ASF under various sensor combinations on K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.3 for ‘Sedan’ and ‘Bus or Truck’ classes. ‘Ove.’ and ‘H.s.’ refer to ‘Overcast’ and ‘Heavy Snow’, respectively. * denotes the sensor unavailability, as shown in Fig. 3. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained for R+L+C. Performance under additional weather conditions (Fog, Rain, and Light Snow) and with other evaluation metric $AP_{BEV}$ is provided in Appendix C.
Detection Performance. As shown in Tab. 1, ASF significantly outperforms SOTA methods for various weather conditions. Compared to the previous SOTA method L4DR [34], ASF shows substantial improvements of 9.7% in $AP_{BEV}$ (i.e., ASF 87.2% vs. L4DR 77.5%) and 20.1% in $AP_{3D}$ (i.e., ASF 73.6% vs. L4DR 53.5%) at IoU=0.5. These improvements are particularly significant for IoU=0.5 (a strict evaluation metric), highlighting ASF’s superior detection performance. The performance advantages are consistent across all weather conditions, with especially notable improvements in challenging sensor degradation scenarios such as sleet (67.5% vs. 46.2% $AP_{3D}$ at IoU=0.5) and heavy snow (66.4% vs. 37.0% $AP_{3D}$ at IoU=0.5).
In addition, we note both ASF (L+R and C+L+R) use the same neural network weights, yet maintain comparable high performance levels. This means that even when camera images become severely degraded, the system gracefully maintains detection performance with the remaining sensors (L+R). This capability is crucial for reliable AD that should be able to cope with unexpected sensor failure. As in Tab. 1, ASF using only R+L achieves 87.0% $AP_{BEV}$ and 72.9% $AP_{3D}$ at IoU=0.5, which is very close to the performance of the full sensor suite (87.2% $AP_{BEV}$ and 73.6% $AP_{3D}$).
Computational Efficiency. The computational efficiency of ASF is another important requirement for real-time AD. As shown in Tab. 2, ASF achieves significant fast processing speed while maintaining low memory requirements. ASF with LiDAR and 4D Radar operates at 20.5 Hz, which is about 4 times (i.e., 4$\times$) faster than 3D-LRF (5.04 Hz) when using the same sensors. Even when all three sensors (C+L+R) are used, ASF maintains 13.5 Hz that exceeds the typical requirement of 10 Hz for AD [8]. This efficiency stems from the proposed CASAP-PN, where cross-attention is applied across sensors along patches rather than across all sensors and all patches as in SCF methods. Additionally, our unified canonical projection eliminates the requirement for computationally expensive positional embedding. The memory footprint of ASF remains remarkably low (1.5-1.6 GB), which is comparable to 3D-LRF (1.2 GB) and much lower than DPFT (4.0 GB).
Exp. $P$ $C_{u}$ $n_{p}$ $n_{h}$ SCL $AP_{3D}^{Sedan}$ $AP_{3D}^{Bus}$
@@
Table 4: Ablation study of ASF. ASF performance for different components and parameters: $P$ (patch size), $C_{u}$ (channel dimension in unified canonical space), $n_{p}$ (number of patches multiplier), $n_{h}$ (number of heads in CASAP-PN) and SCL, using $AP_{3D}$ at IoU=0.3 for both ‘Sedan’ and ‘Bus or Truck’ on the K-Radar benchmark v2.0.
4.3 Addressing Sensor Degradation and Failure
A key advantage of ASF is the robust performance under sensor degradation or failure. As shown in Tab. 3 and Fig. 3, ASF dynamically adapts to different sensor combinations without retraining. Under normal conditions (Fig. 3-(a-i)), ASF effectively utilizes all available sensors with attention weights distributed according to each sensor’s reliability. However, ASF’s true value emerges in challenging scenarios. In adverse weather (Fig. 3-(j-r)), camera and LiDAR measurements are significantly degraded or disappear completely. In these critical situations, ASF automatically redistributes attention toward the more reliable 4D Radar, as evidenced by the predominant blue coloration in the sensor attention maps (SAMs) and corresponding attention percentages. Even with damaged sensors (denoted by * in Tab. 3 and Fig. 3), ASF maintains near-optimal performance. For example, with a damaged camera (C*), C*+L+R shows 77.6% AP3D, which is only 1.7% lower than with fully functional sensors (79.3%). This robustness stems from the unified canonical projection (which creates a common feature space) and the cross-attention mechanism (which implicitly estimates sensor reliability).
4.4 Ablation Studies
Tab. 4 presents ablation studies of key components of ASF. We analyze five factors: patch size ($P$), channel dimension ($C_{u}$), patches multiplier ($n_{p}$), number of attention heads ($n_{h}$), and sensor combination loss (SCL). Comparing experiments (a)-(c), we observe that smaller patch sizes ($P=2$) improve performance by allowing more fine-grained feature extraction. Experiment (d) shows that reducing channel dimension from $C_{u}=512$ to $C_{u}=256$ while increasing the patches multiplier ($n_{p}$) from 4 to 8 maintains or improves the performance, suggesting a better balance between channel depth and spatial representation. Comparing experiments (d) and (f), we find that the number of attention heads impacts classes differently - more heads ($n_{h}=16$) significantly benefit the ’Bus or Truck’ class while showing minimal effect on the ’Sedan’ class. Experiments (e) and (g) demonstrate that incorporating SCL consistently improves performance across configurations, confirming that explicitly training for multiple sensor combinations enhances robustness during inference. The optimal configuration (g) uses $P\!=\!2$, $C_{u}\!=\!256$, $n_{p}\!=\!8$, $n_{h}\!=\!16$, and includes SCL.

### v2

@@ 3.2 Availability-aware Sensor Fusion (ASF)
Figure 3: Qualitative results of ASF for various sensor combinations. We show results for normal and adverse weather conditions in (a-i) and (j-r), respectively, where employed sensors are noted in the top-left corner (C: Camera, L: LiDAR, R: 4D Radar, C*: damaged camera). Each subplot visualizes front-view camera image, LiDAR point cloud, 4D Radar tensor, and a sensor attention map (SAM) showing attention score distribution from cross-attention in CASAP. In the SAMs, red, green, and blue represent attention scores for Camera, LiDAR, and 4D Radar, respectively. For example, a predominantly blue SAM indicates that 4D Radar receives the highest attention scores, meaning that 4D Radar is used primarily for detection in the scene. The bottom-left corner of each subplot shows the proportion of attention scores in colored percentages (C/L/R[%]). Note that predictions are visualized on all sensor data, even when a sensor is not employed for detection (e.g., predictions from L+R are also visualized on the camera image).
Table 1: Performance comparison of 3D object detection on K-Radar (kradar) benchmark v1.0. C, L, and R represent Camera, LiDAR, and 4D Radar, respectively. The bold and underlined values indicate the best and the second-best, respectively. Note that the two ASF models for sensors L+R and C+L+R share the same neural network weights trained with C+L+R, which means that ASF for L+R represents a scenario where camera becomes unavailable. Nor., Ove., Sle., L.s., and H.s. denote ‘Normal’, ‘Overcast’, ‘Sleet’, ‘Light snow’, and ‘Heavy snow’, respectively.
Methods Sensors IoU Metric Total Nor. Ove. Fog Rain Sle. L.s. H.s.
RTNH (kradar) R 0.3 BEV 41.1 41.0 44.6 45.4 32.9 50.6 81.5 56.3
3D 37.4 37.6 42.0 41.2 29.2 49.1 63.9 43.1
0.5 BEV 36.0 35.8 41.9 44.8 30.2 34.5 63.9 55.1
3D 14.1 19.7 20.5 15.9 13.0 13.5 21.0 6.36
RTNH (kradar) L 0.3 BEV 76.5 76.5 88.2 86.3 77.3 55.3 81.1 59.5
3D 72.7 73.1 76.5 84.8 64.5 53.4 80.3 52.9
0.5 BEV 66.3 65.4 87.4 83.8 73.7 48.8 78.5 48.1
3D 37.8 39.8 46.3 59.8 28.2 31.4 50.7 24.6
3D-LRF (lrf) L+R 0.3 BEV 84.0 83.7 89.2 95.4 78.3 60.7 88.9 74.9
3D 74.8 81.2 87.2 86.1 73.8 49.5 87.9 67.2
0.5 BEV 73.6 72.3 88.4 86.6 76.6 47.5 79.6 64.1
3D 45.2 45.3 55.8 51.8 38.3 23.4 60.2 36.9
L4DR (l4dr) L+R 0.3 BEV 79.5 86.0 89.6 89.9 81.1 62.3 89.1 61.3
3D 78.0 77.7 80.0 88.6 79.2 60.1 78.9 51.9
0.5 BEV 77.5 76.8 88.6 89.7 78.2 59.3 80.9 53.8
3D 53.5 53.0 64.1 73.2 53.8 46.2 52.4 37.0
ASF (Proposed) L+R 0.3 BEV 88.6 88.1 90.3 99.0 89.1 80.4 89.4 78.7
3D 87.3 86.6 89.8 90.7 88.6 80.0 88.8 77.5
0.5 BEV 87.0 86.2 90.2 90.8 88.8 78.2 88.6 71.0
3D 72.9 64.6 86.6 79.6 73.4 70.0 77.6 66.7
C+L+R 0.3 BEV 88.6 88.2 90.2 98.9 89.0 80.4 89.2 78.4
3D 87.4 87.0 90.1 90.7 88.2 80.0 88.6 77.4
0.5 BEV 87.2 86.7 90.1 90.8 88.7 78.3 88.3 70.9
3D 73.6 71.8 87.0 79.4 73.0 67.5 78.0 66.4
Table 2: Comparison of VRAM and FPS evaluated on the K-Radar benchmark v1.0. The bold and underlined values indicate the best and the second-best, respectively. The unit of VRAM and FPS are GB and Hz, respectively.
Methods Sensors VRAM FPS
3D-LRF (lrf) L+R 1.2 5.04
DPFT (dpft) C+R 4.0 11.5
ASF L+R 1.5 20.5
(Proposed) C+L+R 1.6 13.5
@@
where $C_{q}$ is the quotient of $C_{u}/(P_{H}\!\times\!P_{W})$ as we design $C_{u}\!=\!P_{H}\!\times\!P_{W}\!\times\!C_{q}$. Since the resulting channel dimension $C_{q}$ may be insufficient for containing feature representation due to reduced channel dimensions after reshaping, in the implementation, we increase the number of patches by a factor of $n_{p}$ (i.e., $N_{p}\!\!=\!\!N_{H}\!\times\!N_{W}\!\!\rightarrow\!\!N_{p}\!=\!n_{p}\!\!\times\!\!N_{H}\!\!\times\!\!N_{W}$). Consequently, the channel dimension of $\mathbf{FM}_{\mathrm{fused}}$ increases from $C_{q}$ to $n_{p}\!\times\!C_{q}$, and impact of this modification is evaluated in subsection 4.4.
@@ 3.3 Sensor Combination Loss (SCL)
where $\mathcal{SC}$ represents the set of $7$ possible sensor combinations ($S_{C}$-only, $S_{L}$-only, $S_{R}$-only, $S_{L}\!+\!S_{R}$, $S_{C}\!+\!S_{R}$, $S_{C}\!+\!S_{L}$, and $S_{C}\!+\!S_{L}\!+\!S_{R}$), where $S_{C}$, $S_{L}$, and $S_{R}$ denote camera, LiDAR, and 4D Radar, respectively. $\mathcal{L}_{cls,s}$ and $\mathcal{L}_{reg,s}$ are the classification and regression losses for each sensor combination. SCL explicitly prepares for the potential sensor unavailability by optimizing across all sensor combinations in the training, enabling the model to recognize that available sensors perform better than others (e.g., 4D Radar outperform camera in adverse weather). As demonstrated in Tab. 4, SCL enhances the performance of the proposed ASF method when compared to ASF without SCL.
@@ 4.1 Experimental Setup
Dataset and Metrics. K-Radar (kradar) is a large-scale autonomous driving dataset with a broad range of conditions including time (day, night), weather (normal, rain, fog, snow, sleet), road types (urban, highway, mountain), and sensors (4D Radar, LiDAR, camera, GPS). Notably, K-Radar is the only dataset with data captured in adverse weather conditions.
For comparison with SOTA methods, we utilize two K-Radar benchmark variants. Benchmark v1.0 (kradar, lrf, l4dr) focuses on the ‘Sedan’ class within a driving corridor region of [0m, 72m] × [-6.4m, 6.4m] × [-2m, 6m] (X×Y×Z). For ablation studies and qualitative analysis, we use benchmark v2.0, which covers a wider area [0m, 72m] × [-16m, 16m] × [-2m, 7.6m] and includes both ‘Sedan’ and ‘Bus or Truck’ classes. We evaluate using $AP_{3D}$ and $AP_{BEV}$ at IoU thresholds of 0.3 and 0.5, while also reporting VRAM usage and FPS based on the same hardware setup.
Table 3: Performance comparison of ASF under various sensor combinations on K-Radar (kradar) benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.3 for ‘Sedan’ and ‘Bus or Truck’ classes. ‘Nor.’, ‘Ove.’, ’Sle.‘ and ‘H.s.’ refer to ‘Normal’, ‘Overcast’, ‘Sleet’, and ‘Heavy snow’, respectively. * denotes the sensor unavailability, as shown in Fig. 3. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained for R+L+C. Performance under additional weather conditions (Fog, Rain, and Light Snow) and with other evaluation metric $AP_{BEV}$ is provided in Appendix D.
Class Sensors Total Nor. Ove. Sle. H.s.
@@
Implementation Details. We implement the ASF on a single RTX3090 GPU with 24GB VRAM. ASF is trained for 11 epochs using AdamW (adamw) optimizer with a learning rate $0.001$ and a batch size $2$. The voxel size for the fused FM is set to 0.40.4m, consistent with (kradar).
4.2 Comparison of ASF to SOTA Methods
Following the benchmark v1.0 of the K-Radar (kradar), we compare the proposed ASF with SOTA methods including 3D-LRF (lrf) and L4DR (l4dr), and we use RTNH (kradar) for single-sensor performance. In addition to the detection performance comparison with DCF methods (lrf, l4dr), we evaluate computational efficiency against DPFT (dpft) which is the only open-sourced SCF method available for K-Radar.
Detection Performance. As shown in Tab. 1, ASF significantly outperforms SOTA methods across various weather conditions. Compared to previous SOTA L4DR (l4dr), ASF achieves substantial improvements of 9.7% in $AP_{BEV}$ (87.2% vs. 77.5%) and 20.1% in $AP_{3D}$ (73.6% vs. 53.5%) at IoU=0.5. These improvements are particularly remarkable in challenging conditions like sleet (67.5% vs. 46.2% $AP_{3D}$) and heavy snow (66.4% vs. 37.0% $AP_{3D}$). Notably, both ASF configurations (L+R and C+L+R) use identical neural network weights yet maintain comparable performance, demonstrating the system’s ability to gracefully handle sensor degradation. Even with only LiDAR and Radar, ASF achieves 87.0% $AP_{BEV}$ and 72.9% $AP_{3D}$ at IoU=0.5, nearly matching the full sensor suite’s performance.
Computational Efficiency. ASF achieves exceptional computational efficiency for real-time autonomous driving applications. As demonstrated in Tab. 2, ASF with LiDAR and 4D Radar processes at 20.5 Hz, approximately 4× faster than 3D-LRF (5.04 Hz) using identical sensors. Even with all three sensors, ASF maintains 13.5 Hz, exceeding the 10 Hz threshold typically required for autonomous driving systems (adsystem_survey). This efficiency results from our CASAP, which applies cross-attention across sensors along patches rather than across all sensors and patches. Furthermore, ASF maintains a compact memory footprint (1.5-1.6 GB), comparable to 3D-LRF (1.2 GB) and substantially lower than DPFT (4.0 GB).
4.3 Addressing Sensor Degradation and Failure
A key advantage of ASF is the robust performance under sensor degradation or failure. As shown in Tab. 3 and Fig. 3, ASF dynamically adapts to different sensor combinations without retraining. Under normal conditions (Fig. 3-(a-i)), ASF effectively utilizes all available sensors with attention weights distributed according to each sensor’s reliability. However, ASF’s true value emerges in challenging scenarios. In adverse weather (Fig. 3-(j-r)), camera and LiDAR measurements are significantly degraded or disappear completely. In these critical situations, ASF automatically redistributes attention toward the more reliable 4D Radar, as evidenced by the predominant blue coloration in the sensor attention maps (SAMs) and corresponding attention percentages. Even with damaged sensors (denoted by * in Tab. 3 and Fig. 3), ASF maintains near-optimal performance. For example, with a damaged camera (C*), C*+L+R shows 77.6% AP3D, which is only 1.7% lower than with fully sensors (79.3%). This robustness stems from the unified canonical projection (which creates a common feature space) and the cross-attention mechanism (which estimates sensor reliability).
4.4 Ablation Studies
Table 4: Ablation study of ASF. ASF performance for different components and parameters: $P$ (patch size), $C_{u}$ (channel dimension in unified canonical space), $n_{p}$ (number of patches multiplier), $n_{h}$ (number of heads in CASAP) and SCL, using $AP_{3D}$ at IoU=0.3 for both ‘Sedan’ and ‘Bus or Truck’ on the K-Radar benchmark v2.0.
Exp. $P$ $C_{u}$ $n_{p}$ $n_{h}$ SCL Sedan Bus
@@ 4.4 Ablation Studies
Tab. 4 presents ablation studies of key ASF components, analyzing five factors: patch size ($P$), channel dimension ($C_{u}$), patches multiplier ($n_{p}$), number of attention heads ($n_{h}$), and sensor combination loss (SCL). Our findings reveal that smaller patch sizes (P=2P\!\!=\!\!2) improve performance through finer feature extraction, while balancing reduced channel dimension ($C_{u}\!\!=\!\!256$) with increased patches multiplier ($n_{p}\!\!=\!\!8$) maintains or enhances results; furthermore, increasing attention heads ($n_{h}\!=\!16$) benefits ‘Bus or Truck’ detection, and incorporating SCL consistently improves performance across configurations by enhancing robustness to varying sensor availability. The optimal configuration combines P=2P\!\!=\!\!2, $C_{u}\!\!=\!\!256$, $n_{p}\!\!=\!\!8$, $n_{h}\!\!=\!\!16$ with SCL.